You have taken a long position in a call option on IBM common stock. The option has an exercise price of $148 and IBM's stock currently trades at $153. The option premium is $7 per contract. a. How much of the option premium is due to intrinsic value versus time value?

b. What is your net profit on the option if IBM’s stock price increases to $163 at expiration of the option and you exercise the option? c. What is your net profit if IBM’s stock price decreases to $143?

How much of the option premium is due to intrinsic value versus time value?

Option Premium
Intrinsic value
Time value

What is your net profit on the option if IBM’s stock price increases to $163 at expiration of the option and you exercise the option and if IBM’s stock price decreases to $143? (Negative amounts should be indicated by a minus sign.)

b. Net profit per share
c. Net profit per share
Answer #1

Exercise price = 148

Current Price = 153

option Premium = 7

a. intrinsic value = current price = exercise price = 153-148 =5

time value = option price - intrinsic value = 7-5 = 2

b. price at expiration = 163

Profit = price at expiration - exercise price - option premium = 163-148-7=8

c. if IBM price falls to 143 , it is below exercise price. Option wont be exercised since market price is less than exercise price. So entire option premium is wasted and loss is option premium = -7
Answer #2

Ans A:

Current Stock Price = $153

Strike Price = $148

So intrinsic value = Current stock price - Strike price

= $153 - $148 = $5

and time value = Option premium - Intrinsic value

= $7 -$5 = $2

So option premium due to Intrinsic value = $5 and due to time value = $2

Ans B

Given Current Market price = $163

and strike price = $148

Of course you would want to exercise the option to buy the stock cheaper than the market price

but for that you will have to pay option premium along with the strike price

so price you pay = $148+ $7 =$$155

and if you sell this stock at the current price level you will make = $163-$155 = $8

Ans C

Given Current Stock price: $143

Now since the market price is lower than the strike price you would not want to exercise the option instead you will purchase the stock from the market at $143.

But you have to pay the premium no matter you exercise the option or not. So you won't make any profit in this situation but lose option premium of $7